# Eliminating Catastrophic Overfitting Via Abnormal Adversarial Examples Regularization

### v1

@@ 1 Introduction
In recent years, deep neural networks (DNNs) have demonstrated impressive performance in various decision-critical domains, such as autonomous driving litman2017autonomous ; grigorescu2020survey , face recognition sharif2016accessorize ; balaban2015deep and medical imaging diagnosis erickson2017machine . However, DNNs were found to be vulnerable to adversarial examples szegedy2013intriguing ; fernandes2020automatic ; huang2023robust . Although these adversarial perturbations are imperceptible to human eyes, they can lead to a completely different prediction in DNNs. To this end, many adversarial defence strategies have been proposed, such as pre-processing techniques guo2017countering , detection algorithms metzen2017detecting , verification and provable defence katz2017reluplex ; zhang2022rethinking , and adversarial training (AT) goodfellow2014explaining ; madry2017towards ; zhang2019theoretically . Among them, AT is considered to be the most effective method against adversarial attacks athalye2018obfuscated ; croce2022evaluating .
Despite the notable progress in improving model robustness, the standard multi-step AT significantly increases the computational overhead due to the iterative steps of forward and backward propagation madry2017towards ; wu2020adversarial ; yu2022robust ; yu2022understanding ; zhou2023phase ; zhou2022improving . In light of this, several works have attempted to use single-step adversarial training (SSAT) as a more efficient alternative to achieve robustness. Unfortunately, a
Figure 1: The test accuracy of RS-FGSM wong2020fast (red line) and RS-AAER (green line) with 16/255 noise magnitude. The dashed and solid lines denote natural and robust (PGD-7-1) accuracy, respectively. The dashed black line corresponds to the 9th epoch, which is the point that RS-FGSM occurs CO.
serious problem catastrophic overfitting (CO) has been identified in SSAT wong2020fast , manifesting as a sharp decline in the model’s robust accuracy against multi-step adversarial attacks, plummeting from a peak to nearly 0% within a few iterations, as shown in Figure 1. This intriguing phenomenon has been widely investigated and prompted numerous efforts to resolve it. Recently, Kim kim2021understanding pointed out that the SSAT-trained classifiers are typically accompanied by highly distorted decision boundaries, which will lead to the model manifestation as CO. However, the underlying process of the classifier’s gradual distortion, as well as the factor inducing rapid distortion, remains unclear.
@@ 2.1 Adversarial Training
DNNs are known to be vulnerable to adversarial attacks szegedy2013intriguing , and AT has been demonstrated to be the most effective defence method athalye2018obfuscated . AT is generally formulated as a min-max optimization problem madry2017towards ; croce2022evaluating . The inner maximization problem tries to generate the strongest adversarial examples to maximize the loss, and the outer minimization problem tries to optimize the network to minimize the loss on adversarial examples, which can be formalized as follows:
@@ 2.2 Catastrophic Overfitting
Since the intriguing phenomenon of CO was identified wong2020fast , there has been a line of work trying to explore and mitigate this problem. wong2020fast  first suggested using a random initialization and early stopping to avoid CO. Furthermore, vivek2020single  empirically showed that using a dynamic dropout schedule can avoid early overfitting to adversarial examples, and de2022make ; zhang2022noise found that incorporating a stronger data augmentation is effective in avoiding CO. Another alternative approach imports partial multi-step AT, for example, wang2021multi  periodically trained the model on natural, single-step and multi-step adversarial examples, and vivek2020plug  built a regularization term by comparing with the multi-step adversarial examples.
However, the above methods have not provided a deeper insight into the essence of CO. Separate works found that CO is closely related to anomalous gradient updates. li2022subspace  constrained the training samples to a carefully extracted subspace to avoid abrupt gradient growth. golgooni2021zerograd  ignored the small gradient adversarial perturbations to mitigate substantial weight updates in the network. huang2022fast  proposed an instance-adaptive SSAT approach where the perturbation size is inversely proportional to the gradient. park2021reliably  leveraged the latent representation of gradients as the adversarial perturbation to compensate for local linearity. sriramanan2021towards  introduced a relaxation term to find more suitable gradient directions by smoothing the loss surface. andriushchenko2020understanding  proposed a regularization term to avoid the non-linear surfaces around the samples. More recently, kim2021understanding  introduced a new perspective that CO is a manifestation of highly distorted decision boundaries. Accordingly, they proposed to reduce the perturbation size for the already misclassified adversarial examples.
@@ 3.1 Definition and Counting of AAE
Adversarial training employs the most adversarial data to reduce the sensitivity of the network’s outputs w.r.t. adversarial perturbation of the natural data. Consequently, the inner maximization process is expected to generate effective adversarial examples that maximize the classification loss. As empirically demonstrated by kim2021understanding , the decision boundaries of the classifier become highly distorted after the occurrence of CO. In this study, we find that after adding the adversarial perturbation generated by the distorted classifier, the loss of certain training samples unexpectedly decreases. This particular behaviour is illustrated in Figure 2, we can observe that the NAEs (blue) can either lead to the model misclassifications or position themselves closer to the decision boundary after the inner maximization process. In contrast, the AAEs (red) will be located further away from the decision boundary and fail to mislead the classifier after adding the perturbation generated by the distorted classifier. Therefore, we introduce the following formula to define AAEs:
@@
Next, we observe the variation in the number of AAEs throughout the model training process, and the corresponding statistical results are presented in Figure 3 (left). It can be observed that before the occurrence of CO, a small number of AAEs already existed, indicating the presence of slight initial distortion in the classifier. To further validate this point, we visualize the loss surface of both AAEs and NAEs using the method proposed by li2018visualizing as shown in Figure 4 (left and middle). It’s evident that before CO, the classifier showcases a more nonlinear loss surface around AAEs in comparison to NAEs. This empirical observation strongly suggests that the generation of AAEs is directly influenced by the distorted classifier.
@@ 4.1 Experiment Settings
Baselines.  We compare our method with other SSAT methods, including RS-FGSM wong2020fast , FreeAT shafahi2019adversarial , N-FGSM de2022make , Grad Align andriushchenko2020understanding , ZeroGrad and MultiGrad golgooni2021zerograd . We also compare our method with multi-step AT, PGD-2 and PGD-10 (PGD-20 with 32/255 noise magnitude) madry2017towards , providing a reference for the ideal performance. The results of other competing baselines, including GAT sriramanan2020guided , NuAT sriramanan2021towards , PGI-FGSM jia2022prior , SDI-FGSM jia2022boosting and Kim kim2021understanding , can be found in Appendix F. We report both the natural and robust accuracy results of the final model, which are obtained without early stopping and using the hyperparameters provided in the official repository. Please note that for FreeAT, we did not use the subset of training samples to keep the same training epochs across different methods.
Attack Methods.  To report the robust accuracy of models, we attack these methods using the standard PGD adversarial attack with $\alpha=\epsilon/4$ step size, 50 attack steps and 10 restarts. We also evaluate our methods on Auto Attack croce2020reliable as shown in Appendix C.
Datasets and Model Architectures.  We evaluate our method on several benchmark datasets, including Cifar-10/100 krizhevsky2009learning , SVHN netzer2011reading , Tiny-ImageNet netzer2011reading and Imagenet-100 deng2009imagenet . The standard data augmentation random cropping and horizontal flipping are applied for these datasets. The settings and results on SVHN, Tiny-ImageNet and Imagenet-100 are provided in Appendix E. We use the PreactResNet-18 he2016identity and WideResNet-34 zagoruyko2016wide architectures on these datasets to evaluate results. The results of WideResNet-34 can be found in Appendix D.
Setup for Our Proposed Method.  In this work, we use the SGD optimizer with a momentum of 0.9, weight decay of 5 × $10^{-4}$ and $L_{\infty}$ as the threat model. For the learning rate schedule, we use the cyclical learning rate schedule smith2017cyclical with 30 epochs, which reaches its maximum learning rate (0.2) when half of the epochs (15) are passed. The results obtained by the long training schedule can be found in Appendix G. We superimpose our method on two baseline methods: RS-FGSM and N-FGSM, both of which use the vanilla min-max process. For RS-AAER, we follow the settings of wong2020fast that set step size $\alpha=1.25\cdot\epsilon$ and random initialization $\eta=\text{Uniform}(-\epsilon,\epsilon)$. In accordance with the N-FGSM setting suggested by de2022make , we set $\alpha=1.0\cdot\epsilon$ and $\eta=\text{Uniform}(-2\cdot\epsilon,2\cdot\epsilon)$ for N-AAER. The hyperparameter settings for Cifar-10/100 are summarized in the Table 1.
@@ Appendix A Vanilla-AAER
To further validate the effectiveness of our method, we implement Vanilla-AAER to prevent CO. The Vanilla-AAER method follows the settings of Vanilla-FGSM goodfellow2014explaining , which does not use random initialization and sets the step size as $\alpha=1.0\cdot\epsilon$. The Vanilla-AAER hyperparameters setting is shown in Table 5.
@@
Based on the results presented in Table 6, we can observe that Vanilla-AAER achieves comparable or even superior robustness compared to RS-AAER. This outcome may be attributed to the fact that Vanilla-FGSM has a higher expectation of adversarial perturbation, as demonstrated in prior work andriushchenko2020understanding . However, the absence of random initialization in Vanilla-AAER may reduce the diversity of adversarial examples, potentially impacting the natural accuracy of the model. Nonetheless, the most significant finding is that Vanilla-AAER effectively eliminates CO across various noise magnitudes, which cannot be accomplished by Vanilla-FGSM.
@@ Appendix B Impacts of Regularization Term
To showcase the distinct effectiveness of our method in eliminating CO, we conducted a comparison with other regularization methods used in multi-step AT, such as TRADES zhang2019theoretically and ALP kannan2018adversarial . In order to ensure a fair comparison, we set the iteration times for TRADES and ALP as 1, and the step size as $\alpha=1.25\cdot\epsilon$ and $\alpha=1.0\cdot\epsilon$ for superimposing RS-FGSM and N-FGSM, respectively.
@@ Appendix C Evaluation Based on Auto Attack
Auto Attack croce2020reliable is regarded as the most reliable robustness evaluation to date. It is an ensemble of complementary attacks, consisting of three white-box attacks (APGD-CE, APGD-DLR, and FAB) and a black-box attack (Square Attack). In order to avoid the pseudo-robustness brought by gradient masking or gradient obfuscation, we report the Auto Attack results on Cifar10/100 in Table 8 and Table 9.
@@ Appendix E Settings and Results on SVHN, Tiny-ImageNet and Imagenet-100
SVHN Settings and Results.  For experiments on SVHN, we use the cyclical learning rate schedule with 15 epochs that reaches its maximum learning rate (0.05) when 40% (6) epochs are passed. In the meantime, we uniformly increase the step size between 0 and 5 epochs, which follow the settings of de2022make . We show the hyperparameters setting on SVHN in Table 12. In Table 13, we present the performance of AAER on the SVHN dataset, along with the results of the competing baseline taken from de2022make . It is evident that our method can successfully prevent CO and improve robust accuracy across different noise magnitudes. This demonstrates the effectiveness of AAER in enhancing the robustness of models trained on the SVHN dataset.
@@
Table 13: SVHN: Accuracy of different methods and different noise magnitudes using PreActResNet-18 under $L_{\infty}$ threat model. The baseline results are taken from de2022make . The top number is the natural accuracy (%), while the bottom number is the PGD-50-10 accuracy (%). The results are averaged over 3 random seeds and reported with the standard deviation.
@@ Appendix F More Competing Baselines
We also compare the performance of our method with other SSAT methods, including GAT sriramanan2020guided , NuAT sriramanan2021towards , PGI-FGSM jia2022prior , SDI-FGSM jia2022boosting and Kim kim2021understanding . The results for GAT, NuAT and Kim are directly taken from de2022make , while the reported PGI-FGSM and SDI-FGSM results are based on the official code after searching the hyperparameters across different noise magnitudes.
Table 16: CIFAR10: Accuracy of different methods and different noise magnitudes using PreActResNet-18 under $L_{\infty}$ threat model. GAT, NuAT and Kim results are taken from de2022make . The top number is the natural accuracy (%), while the bottom number is the PGD-50-10 accuracy (%). The results are averaged over 3 random seeds and reported with the standard deviation.

### v2

@@ 1 Introduction
In recent years, deep neural networks (DNNs) have demonstrated impressive performance in various decision-critical domains, such as autonomous driving [25, 12], face recognition [31, 3] and medical imaging diagnosis [8]. However, DNNs were found to be vulnerable to adversarial examples [35, 9, 16]. Although these adversarial perturbations are imperceptible to human eyes, they can lead to a completely different prediction in DNNs. To this end, many adversarial defence strategies have been proposed, such as pre-processing techniques [13], detection algorithms [27], verification and provable defence [20, 44], and adversarial training (AT) [11, 26, 46]. Among them, AT is considered to be the most effective method against adversarial attacks [2, 4].
Despite the notable progress in improving model robustness, the standard multi-step AT significantly increases the computational overhead due to the iterative steps of forward and backward propagation [26, 40, 41, 42, 48, 47]. In light of this, several works have attempted to use single-step adversarial training (SSAT) as a more efficient alternative to achieve robustness. Unfortunately, a
Figure 1: The test accuracy of RS-FGSM [39] (red line) and RS-AAER (green line) with 16/255 noise magnitude. The dashed and solid lines denote natural and robust (PGD-7-1) accuracy, respectively. The dashed black line corresponds to the 9th epoch, which is the point that RS-FGSM occurs CO.
serious problem catastrophic overfitting (CO) has been identified in SSAT [39], manifesting as a sharp decline in the model’s robust accuracy against multi-step adversarial attacks, plummeting from a peak to nearly 0% within a few iterations, as shown in Figure 1. This intriguing phenomenon has been widely investigated and prompted numerous efforts to resolve it. Recently, Kim [21] pointed out that the SSAT-trained classifiers are typically accompanied by highly distorted decision boundaries, which will lead to the model manifestation as CO. However, the underlying process of the classifier’s gradual distortion, as well as the factor inducing rapid distortion, remains unclear.
@@ 2.1 Adversarial Training
DNNs are known to be vulnerable to adversarial attacks [35], and AT has been demonstrated to be the most effective defence method [2]. AT is generally formulated as a min-max optimization problem [26, 4]. The inner maximization problem tries to generate the strongest adversarial examples to maximize the loss, and the outer minimization problem tries to optimize the network to minimize the loss on adversarial examples, which can be formalized as follows:
@@ 2.2 Catastrophic Overfitting
Since the intriguing phenomenon of CO was identified [39], there has been a line of work trying to explore and mitigate this problem. [39] first suggested using a random initialization and early stopping to avoid CO. Furthermore, [36] empirically showed that using a dynamic dropout schedule can avoid early overfitting to adversarial examples, and [6, 45] found that incorporating a stronger data augmentation is effective in avoiding CO. Another alternative approach imports partial multi-step AT, for example, [38] periodically trained the model on natural, single-step and multi-step adversarial examples, and [37] built a regularization term by comparing with the multi-step adversarial examples.
However, the above methods have not provided a deeper insight into the essence of CO. Separate works found that CO is closely related to anomalous gradient updates. [24] constrained the training samples to a carefully extracted subspace to avoid abrupt gradient growth. [10] ignored the small gradient adversarial perturbations to mitigate substantial weight updates in the network. [15] proposed an instance-adaptive SSAT approach where the perturbation size is inversely proportional to the gradient. [29] leveraged the latent representation of gradients as the adversarial perturbation to compensate for local linearity. [34] introduced a relaxation term to find more suitable gradient directions by smoothing the loss surface. [1] proposed a regularization term to avoid the non-linear surfaces around the samples. More recently, [21] introduced a new perspective that CO is a manifestation of highly distorted decision boundaries. Accordingly, they proposed to reduce the perturbation size for the already misclassified adversarial examples.
@@ 3.1 Definition and Counting of AAE
Adversarial training employs the most adversarial data to reduce the sensitivity of the network’s outputs w.r.t. adversarial perturbation of the natural data. Consequently, the inner maximization process is expected to generate effective adversarial examples that maximize the classification loss. As empirically demonstrated by [21], the decision boundaries of the classifier become highly distorted after the occurrence of CO. In this study, we find that after adding the adversarial perturbation generated by the distorted classifier, the loss of certain training samples unexpectedly decreases. This particular behaviour is illustrated in Figure 2, we can observe that the NAEs (blue) can either lead to the model misclassifications or position themselves closer to the decision boundary after the inner maximization process. In contrast, the AAEs (red) will be located further away from the decision boundary and fail to mislead the classifier after adding the perturbation generated by the distorted classifier. Therefore, we introduce the following formula to define AAEs:
@@
Next, we observe the variation in the number of AAEs throughout the model training process, and the corresponding statistical results are presented in Figure 3 (left). It can be observed that before the occurrence of CO, a small number of AAEs already existed, indicating the presence of slight initial distortion in the classifier. To further validate this point, we visualize the loss surface of both AAEs and NAEs using the method proposed by [23] as shown in Figure 4 (left and middle). It’s evident that before CO, the classifier showcases a more nonlinear loss surface around AAEs in comparison to NAEs. This empirical observation strongly suggests that the generation of AAEs is directly influenced by the distorted classifier.
@@ 4.1 Experiment Settings
Baselines.  We compare our method with other SSAT methods, including RS-FGSM [39], FreeAT [30], N-FGSM [6], Grad Align [1], ZeroGrad and MultiGrad [10]. We also compare our method with multi-step AT, PGD-2 and PGD-10 (PGD-20 with 32/255 noise magnitude) [26], providing a reference for the ideal performance. The results of other competing baselines, including GAT [33], NuAT [34], PGI-FGSM [17], SDI-FGSM [18] and Kim [21], can be found in Appendix F. We report both the natural and robust accuracy results of the final model, which are obtained without early stopping and using the hyperparameters provided in the official repository. Please note that for FreeAT, we did not use the subset of training samples to keep the same training epochs across different methods.
Attack Methods.  To report the robust accuracy of models, we attack these methods using the standard PGD adversarial attack with $\alpha=\epsilon/4$ step size, 50 attack steps and 10 restarts. We also evaluate our methods on Auto Attack [5] as shown in Appendix C.
Datasets and Model Architectures.  We evaluate our method on several benchmark datasets, including Cifar-10/100 [22], SVHN [28], Tiny-ImageNet [28] and Imagenet-100 [7]. The standard data augmentation random cropping and horizontal flipping are applied for these datasets. The settings and results on SVHN, Tiny-ImageNet and Imagenet-100 are provided in Appendix E. We use the PreactResNet-18 [14] and WideResNet-34 [43] architectures on these datasets to evaluate results. The results of WideResNet-34 can be found in Appendix D.
Setup for Our Proposed Method.  In this work, we use the SGD optimizer with a momentum of 0.9, weight decay of 5 × $10^{-4}$ and $L_{\infty}$ as the threat model. For the learning rate schedule, we use the cyclical learning rate schedule [32] with 30 epochs, which reaches its maximum learning rate (0.2) when half of the epochs (15) are passed. The results obtained by the long training schedule can be found in Appendix G. We superimpose our method on two baseline methods: RS-FGSM and N-FGSM, both of which use the vanilla min-max process. For RS-AAER, we follow the settings of [39] that set step size $\alpha=1.25\cdot\epsilon$ and random initialization $\eta=\text{Uniform}(-\epsilon,\epsilon)$. In accordance with the N-FGSM setting suggested by [6], we set $\alpha=1.0\cdot\epsilon$ and $\eta=\text{Uniform}(-2\cdot\epsilon,2\cdot\epsilon)$ for N-AAER. The hyperparameter settings for Cifar-10/100 are summarized in the Table 1.
@@ Appendix A Vanilla-AAER
To further validate the effectiveness of our method, we implement Vanilla-AAER to prevent CO. The Vanilla-AAER method follows the settings of Vanilla-FGSM [11], which does not use random initialization and sets the step size as $\alpha=1.0\cdot\epsilon$. The Vanilla-AAER hyperparameters setting is shown in Table 5.
@@
Based on the results presented in Table 6, we can observe that Vanilla-AAER achieves comparable or even superior robustness compared to RS-AAER. This outcome may be attributed to the fact that Vanilla-FGSM has a higher expectation of adversarial perturbation, as demonstrated in prior work [1]. However, the absence of random initialization in Vanilla-AAER may reduce the diversity of adversarial examples, potentially impacting the natural accuracy of the model. Nonetheless, the most significant finding is that Vanilla-AAER effectively eliminates CO across various noise magnitudes, which cannot be accomplished by Vanilla-FGSM.
@@ Appendix B Impacts of Regularization Term
To showcase the distinct effectiveness of our method in eliminating CO, we conducted a comparison with other regularization methods used in multi-step AT, such as TRADES [46] and ALP [19]. In order to ensure a fair comparison, we set the iteration times for TRADES and ALP as 1, and the step size as $\alpha=1.25\cdot\epsilon$ and $\alpha=1.0\cdot\epsilon$ for superimposing RS-FGSM and N-FGSM, respectively.
@@ Appendix C Evaluation Based on Auto Attack
Auto Attack [5] is regarded as the most reliable robustness evaluation to date. It is an ensemble of complementary attacks, consisting of three white-box attacks (APGD-CE, APGD-DLR, and FAB) and a black-box attack (Square Attack). In order to avoid the pseudo-robustness brought by gradient masking or gradient obfuscation, we report the Auto Attack results on Cifar10/100 in Table 8 and Table 9.
@@ Appendix E Settings and Results on SVHN, Tiny-ImageNet and Imagenet-100
SVHN Settings and Results.  For experiments on SVHN, we use the cyclical learning rate schedule with 15 epochs that reaches its maximum learning rate (0.05) when 40% (6) epochs are passed. In the meantime, we uniformly increase the step size between 0 and 5 epochs, which follow the settings of [6]. We show the hyperparameters setting on SVHN in Table 12. In Table 13, we present the performance of AAER on the SVHN dataset, along with the results of the competing baseline taken from [6]. It is evident that our method can successfully prevent CO and improve robust accuracy across different noise magnitudes. This demonstrates the effectiveness of AAER in enhancing the robustness of models trained on the SVHN dataset.
@@
Table 13: SVHN: Accuracy of different methods and different noise magnitudes using PreActResNet-18 under $L_{\infty}$ threat model. The baseline results are taken from [6]. The top number is the natural accuracy (%), while the bottom number is the PGD-50-10 accuracy (%). The results are averaged over 3 random seeds and reported with the standard deviation.
@@ Appendix F More Competing Baselines
We also compare the performance of our method with other SSAT methods, including GAT [33], NuAT [34], PGI-FGSM [17], SDI-FGSM [18] and Kim [21]. The results for GAT, NuAT and Kim are directly taken from [6], while the reported PGI-FGSM and SDI-FGSM results are based on the official code after searching the hyperparameters across different noise magnitudes.
Table 16: CIFAR10: Accuracy of different methods and different noise magnitudes using PreActResNet-18 under $L_{\infty}$ threat model. GAT, NuAT and Kim results are taken from [6]. The top number is the natural accuracy (%), while the bottom number is the PGD-50-10 accuracy (%). The results are averaged over 3 random seeds and reported with the standard deviation.